OpenStudy (anonymous):

Find the range of f(x)=sqrt(9-x^2) algebraically.

OpenStudy (anonymous):

The domain is (-infinity, 3] But I only know how to find the range by graphing, not algebraically.

OpenStudy (amistre64):

start with the domain and work it into the function

OpenStudy (anonymous):

How would I do that?

OpenStudy (amistre64):

-inf < x <= 3 since y = f(x), convert x into the f(x) square, negate, sqrt

OpenStudy (anonymous):

What do you mean convert x into the f(x)?

myininaya (myininaya):

If I were your cat, I would lick your domain off your paper.

OpenStudy (amistre64):

start with the domain, this < x < that then convert x into f(x) by the appropriate transformations

OpenStudy (amistre64):

the range is some < f(x) < thing

myininaya (myininaya):

unless you told me you wanted to include some complex numbers and not others

OpenStudy (amistre64):

is it the notation f(x) that you dont understand? it simply means a function of x, which you already have defined

OpenStudy (anonymous):

So would it be -inf < f(x) < 3?

OpenStudy (amistre64):

yeah, clean up your domain 9-x^2 >= 0 9 >= x^2 x <= 3 and x >= -3

OpenStudy (anonymous):

My teacher makes us put it in interval notation, sorry.

OpenStudy (amistre64):

im trying to rush something in my head ....

OpenStudy (amistre64):

x=0 works, but any x^2 bigger then 9 goes negative ...

OpenStudy (amistre64):

so -3 to 3 is the domain, agreed?

OpenStudy (anonymous):

Wait, so would the domain be (-inf, -3] U [3, inf)?

OpenStudy (anonymous):

Yes, what is that in interval notation so I can write that down?

OpenStudy (amistre64):

9 - (inf)^2 = -inf

OpenStudy (amistre64):

anything bigger than 9 goes negatice

OpenStudy (anonymous):

I am so confused lol I have no idea what you are doing.

OpenStudy (amistre64):

can we have a negative number under the radical?

OpenStudy (anonymous):

My teacher never taught us this and I got it wrong on the test so I'm trying to figure out how to do everything I got wrong for future tests. No you cannot or else it will be imaginary.

OpenStudy (amistre64):

then we have a postive number, 9 and a number that is subtracted from it, do you see that when we subtract something greater than 9, from 9, we go negative?

OpenStudy (anonymous):

Yes. I see that.

OpenStudy (amistre64):

then the only values for the domain are from -3 to 3

OpenStudy (amistre64):

D = [-3,3] not -inf to 3

OpenStudy (anonymous):

I figured it out!

OpenStudy (amistre64):

so we have the domain: -3 <= x <= 3 we have a function defined: f(x) = sqrt(9-x^2) the range is defined for some interval: c <= f(x) <= d if we take the domain, and work x into f(x) then the interval itslef is worked into the range

OpenStudy (anonymous):

Domain: 9-x^2 >= 0 -x^2 >= 9 x^2 <= 9 x <= +/- 3 -3 < x < 3

OpenStudy (anonymous):

Interval notation would be [-3, 3]

OpenStudy (amistre64):

yes

OpenStudy (anonymous):

Range: x=sqrt(9-y^2) 9-y^2 >= 0 9 >= y^2 y <= +/- 3 -3 <= x <= 3 [-3, 3]

OpenStudy (amistre64):

now, we may need to do some mathical trick to make this work smoothly would you agree that: -3 <= x <= 3 is that same as |x| <= 3 ??

OpenStudy (amistre64):

that is not the range

OpenStudy (anonymous):

Ugh. I got half of it now though lol.

OpenStudy (anonymous):

I would agree that |x| <= 3.

OpenStudy (amistre64):

lets transform x into f(x) if we try this: -3 <= x <= 3, square it all 9 <= x^2 <= 9, negate it all (and signs flip when we multiply by -1) -9 >= -x^2 >= -9, add 9 0 >= 9-x^2 >= 0, and sqrt 0 >= sqrt(9-x^2) >= 0 which doesnt really make a lot of sense in this form, which is why writing the domain as: |x| <= 3 might work better

OpenStudy (amistre64):

|x| <= 3 (|x|)^2 <= 9, but (|x|)^2 is equal to x^2 x^2 <= 9, negate and flip the sign -9 <= -x^2, add 9 0 <= 9-x^2, and sqrt 0 <= sqrt(9-x^2) i cant seem to algebra the top at the moment, its max is 3 0 <= sqrt(9-x^2) <= 3

OpenStudy (amistre64):

y = sqrt(9-x^2) y^2 = 9-x^2 x^2 + y^2 = 9 is a circle of radius 3 centered at the origin which gives us a max of +3

OpenStudy (anonymous):

How did you find the max? Is it because that is the square root of 9?

OpenStudy (amistre64):

i found it by memory, the equation is the top half of a circle of radius 3

OpenStudy (anonymous):

So x^2 + y^2 makes a circle?

OpenStudy (amistre64):

yes

OpenStudy (anonymous):

So 0 <= y <= 3 Interval notation making it [0, 3]

OpenStudy (amistre64):

yes, but how to find the 3 eludes me algebraically

OpenStudy (anonymous):

Well, I'm more knowledgeable about this subject than I was before, so it's a start. Thank you anyways!

OpenStudy (amistre64):

good luck with it :)